# Learnable Sparsity for Vision Generative Models

### v2

@@ 1 Introduction
Recently, generative models, such as diffusion models and flow matching models, have made remarkable progress in various vision tasks, including text-to-image generation (brooks2023instructpix2pix; rombach2022high; ramesh2022hierarchical; nichol2022glide), image inpainting (saharia2022palette; lugmayr2022repaint; corneanu2024latentpaint), super-resolution (saharia2022image; li2022srdiff), and video generation (ho2022imagen; luo2023videofusion; singermake). This progress has been marked by the evolution of generative models toward increasingly larger sizes, transitioning from the U-Net-based Stable Diffusion 1 (SD1) (rombach2022high) to the larger SDXL (podell2024sdxl), the transformer-based Stable Diffusion 3 (SD3), and more recently, the FLUX model (esser2024scaling; blackforestlabs2024). The most recent FLUX model, with 12 billion parameters, is about 13 times larger than the SD2, which was developed just two years ago (blackforestlabs2024). The rapid growth in size has introduced challenges in application: larger models demand larger GPUs and more computation during inference, limit deployment on smaller compute platforms, and substantially increase the carbon footprint.
@@
Due to the potential challenges induced by larger models, prior methods have explored ways to reduce model size and computation, including model distillation (meng2023distillation; huang2024knowledge) and model pruning (fang2023structural). Compared to standard knowledge distillation that usually trains a student model from scratch, pruning is computationally more efficient. However, pruned models with fewer weights often suffer from degraded generative performance, making retraining necessary to restore their capabilities. fang2023structural estimate that diffusion model compression can demand 10% to 20% of the original training cost. Therefore, compressing an SD2 can consume up to $40,000$ GPU hours (sd2cost). Recently, kim2024bk showed that the cost of pruning SD2 can be reduced to around 300 A100 GPU hours with 0.22M calibration data. For pruning larger models, such as SDXL and FLUX, the retraining burden is more substantial.
@@
One primary reason for the extensive retraining required in prior approaches is due to their coarse pruning criteria. Prior pruning approaches for diffusion models use simple heuristics or one-shot pruning strategies (kim2024bk; fang2023structural), which often fail to balance sparsity and performance. To improve the pruning selection criteria, differentiable masking has been explored in classical pruning research (louizos2018learning; wen16learning_structured; Liu_2015_CVPR; deep_compression; yuan2006model) and has recently been adapted to large language models (LLMs) (fang2024maskllm). However, several unique characteristics of vision generative models make the application of differentiable masking difficult. For instance, vision generative models are Markovian (songdenoising; songscore; lipmanflow), where each generation step only relies on the previous state. Consequently, a minor change at an intermediate generation step can cause a ripple effect that significantly distorts the final result.
@@
While prior works often require extensive retraining to recover model quality after pruning, we show that this step can be made far more lightweight. By combining pruning with targeted post-pruning adaptation, such as low-rank (LoRA) fine-tuning (hu2022lora) or full-model retraining, our approach enables substantial quality recovery without prohibitive compute. This makes pruning practical even for large-scale diffusion models like SDXL, where retraining can otherwise be resource-intensive. However, end-to-end mask learning results in an extremely long gradient chain across all generation steps, leading to significant memory demands. For instance, performing end-to-end pruning on SDXL requires approximately 1400 GB of VRAM–equivalent to the capacity of $15$ NVIDIA H100 GPUs. To address this, we adopt gradient checkpointing (chen2016training) and introduce time step gradient checkpointing. This approach reduces the VRAM usage for SDXL from 1400 GB to under 30 GB, with a minor overhead of an additional forward pass. An overview of end-to-end pruning and gradient checkpointing is shown in Figure 2.
We show that our method is general and scalable by pruning the most recent diffusion model (SDXL) and flow matching model (FLUX) in Section 6.2. Recent studies have shown that step distillation can greatly speed up inference by reducing generation steps (salimansprogressive; meng2023distillation). However, step-distilled models are difficult to retrain due to their non-smoothness (miao2024tuning; jia2024reward). We show that we can prune step-distilled models in Section 6.2 by pruning FLUX-schnell.
@@ 2 Related Work
Efficient diffusion models: Several works focus on enhancing the efficiency of diffusion models at inference time. DDIM formulates diffusion processes with non-Markovian transformations, reducing the required diffusion steps for high-quality generation (songdenoising). The Latent Diffusion Model adopts a two-stage generation process, performing the diffusion process in latent space and decoding it back to the image domain using a Variational Autoencoder (VAE) (rombach2022high; kingma2014auto). In addition, some prior works accelerate diffusion inference by sharing intermediate variables to reduce redundant computations (ma2024deepcache; ma2024learning; so2023frdiff). However, these methods are often limited to specific architectures and do not generalize well to newer diffusion models incorporating multi-modal attention. Early stopping methods attempt to terminate the diffusion process once satisfactory generation is reached, but they can sacrifice fine-grained details (lyu2022accelerating). Distillation-based approaches train a student model that is either smaller or can generate with fewer diffusion steps (salimansprogressive; meng2023distillation; gu2023boot; hsiao2024plug). Nevertheless, model distillation requires extensive retraining.
@@
Diffusion training efficiency: Beyond inference speedup, research also targets the efficiency of diffusion model training. These efforts range from data-centric acceleration, such as lossless speedup through unbiased dynamic data pruning (qin2024infobatch) or approaches for data-efficient training via sample reweighting (li2025pruning), to architectural and objective acceleration that utilizes masked transformers for faster training (zheng2024fast), patched denoising diffusion models for high-resolution synthesis (ding2023patched), general fast diffusion model designs (wu2023fast), and min-SNR weighting strategies (hang2023efficient). These approaches primarily target training-time efficiency, whereas our work focuses on inference-time structural compression and is complementary to the above techniques.
@@
Model pruning: Pruning reduces model size, thereby lowering both memory requirements for loading the model and computation demands during inference. Many classical pruning approaches learn a mask to prune CNN models (louizos2018learning; wen16learning_structured; Liu_2015_CVPR; deep_compression; yuan2006model; Feng_2015_ICCV). Recent advances in model pruning have predominantly focused on large language models (LLMs), where techniques such as unstructured and semi-structured pruning eliminate connections between neurons, and structured pruning targets neurons, attention heads, and layers (kwon2022a; fang2024maskllm; zhang2024finercut; zhang2024investigating; yu2018nisp; ma2023llm; kwon2022fast; sun2023simple). In contrast, pruning methods for diffusion models remain relatively underexplored. These approaches vary in their granularity: some involve structural pruning, while others focus on unstructured pruning. Some structural approaches (castells2024ld; li2023snapfusion; zhao2024mobilediffusion) perform occlusion-based pruning that calculates an importance score for occluding a part through exhaustive search, but its high complexity limits scalability to complex pruning scenarios with many candidates. Unstructured pruning has also been explored, for example, Efficient Pruning of Text-to-Image Models (EPTI) (ramesh2024efficient) uses criteria such as magnitude and WANDA, reporting strong results without demanding the extensive retraining common in many other pruning approaches. fang2023structural uses gradient information as a proxy for neuron importance. Compared to the occlusion method, it scales to more complex pruning cases. Other approaches (kim2024bk) prune certain redundant blocks and recover the model with feature distillation to further reduce the pruning cost. Still, these methods require excessive retraining to retain the model performance.
@@ 3 Preliminaries
Modules in transformer blocks. Transformer blocks are one of the major building blocks in U-Net diffusion models (rombach2022high; podell2024sdxl). In addition, recent diffusion transformers employ fully transformer-based architectures. This makes transformer blocks a primary target for pruning to improve efficiency. A transformer block consists of multi-head attention (MHA) and feed-forward network (FFN) (vaswani2017attention). The MHA with $h$ attention heads is defined as,
@@
where $W_{1}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{ff}}}$ and $W_{2}\in\mathbb{R}^{d_{\text{ff}}\times d_{\text{model}}}$. $\sigma(\cdot)$ represents an activation function like GELU or GeGLU (hendrycks2016gaussian; Shazeer2020GLUVI).
@@ 5.3 Continuous Relaxation of Discrete Masking
Our pruning objective in Equation 6 involves the calculation of $L_{0}$ norm, which is not differentiable. Hence, we adopt the continuous relaxation of the sparse optimization by applying hard-concrete sampling originally proposed by louizos2018learning. For a scalar continuous masking variable $\hat{\mathcal{M}}\in[0,1]$, the hard-concrete sampling is as follows:
@@ 5.4 Time Step Gradient Checkpointing
To perform the end-to-end pruning shown in Algorithm 1, the backpropagation needs to traverse all generation steps. This necessitates storing all intermediate variables across all steps, leading to a substantial increase in memory usage during mask optimization. We implement time-step gradient checkpointing to address the memory challenge. Traditional gradient checkpointing reduces memory usage by storing selected intermediate outputs and recomputing forward passes between checkpoints (chen2016training; gruslys2016memory). However, traditional checkpointing only stores intermediate values within a single forward pass, thereby limiting its utility for diffusion models, which require multiple model forward passes across time steps. To address this, we design time step gradient checkpointing, which stores intermediate denoised latent $\hat{z}$ after each denoising step. For a diffusion process with $T$ steps, end-to-end backpropagation has a memory complexity of $O(T)$. By implementing time-step gradient checkpointing, we reduce memory complexity to $O(1)$, making it independent of $T$. The runtime complexity remains $O(T)$, with the only added cost being an additional forward pass to recompute activations as needed. We present Algorithm 2 to show how to apply time step gradient checkpointing to calculate the loss gradient w.r.t. the mask control variable $\bm{\lambda}$.
@@ 6.1 Setup
Model: We prune state-of-the-art (SOTA) models: SDXL, FLUX-dev, and FLUX-schnell (rombach2022high; podell2024sdxl; blackforestlabs2024). These models cover diffusion and flow-matching, as well as UNet and DiT architecture. Pruning details: We only use $100$ text prompts from GCC3M (sharma2018conceptual) for pruning, and a computation budget of 10 A100 GPU hours. Baselines: We compare our method with the SOTA diffusion pruning methods, DiffPruning (fang2023structural), BK-SDM (kim2024bk), and FLUX-Lite (flux1-lite). For BK-SDM, we only run on SD-XL, since the method is designed for U-Net models. We also include the per-step loss variant for mask learning as a baseline. Per-step loss uses identical settings as EcoDiff, except the masking loss is changed to a per-step reconstruction loss. Evaluation metrics: We select Fréchet Inception Distance (FID) (heusel2017gans), CLIP score (radford2021learning), and Structural Similarity Index Measure (SSIM) (wang2004image). We evaluate under $5,000$ MS COCO and Flickr 30k data samples respectively (lin2014microsoft; young2014image). All experiments use a single NVIDIA A100 GPU with 80GB VRAM. More details in Appendix B.
@@ 6.2 Main Results
SDXL is one of the SOTA diffusion models that employs a U-Net architecture with 2.62.6B parameters (podell2024sdxl). We compare the pruning results with BK-SDM, DiffPruning, and our per-step loss variant. From Table 1 and Figure 3, EcoDiff outperforms baseline methods under the same computation budget. We also observe low SSIM scores for all our pruned models compared to conventional pruning methods (fang2023structural), with all values below 0.65. This suggests that EcoDiff prioritizes semantic integrity (high FID/CLIP) over pixel-level mimicry (low SSIM). The low SSIM is attributed to subtle, perceptually acceptable shifts in generated textures or fine-detail placement, which are heavily penalized by pixel-level metrics but do not detract from the overall image quality or semantic adherence to the prompt. As shown in Figure 3, the semantic fidelity and fine-grained quality of our generated images are noticeably superior to those of the baseline methods. We explain more in Section 6.5 and provide results on SD2 in Appendix D.
@@
Recently, flow-matching has gained attention due to its superior performance. Currently, the best-performing vision generative model is FLUX-Dev, a flow DiT model with 12B parameters (blackforestlabs2024). Due to the DiT architecture, BK-SDM is not applicable to FLUX. Hence, we compare with DiffPruning and FLUX-Lite, a community-pruned model leveraging redundancy of middle layers. According to Table 1, both EcoDiff and FLUX-Lite preserve the model performance. However, the FLUX-Lite 8B model was derived from a large-scale, post-hoc retraining effort, requiring 1120 hours on H200 GPUs to achieve 33% sparsity (huggingface_flux16), we highlight this comparison to demonstrate that EcoDiff achieves comparable quality and structural sparsity (20%) using only 10 A100 GPU hours for the mask learning stage, positioning our method as a vastly more efficient and accessible alternative to resource-intensive retraining campaigns.
@@
FLUX-Schnell is the step-distilled model from FLUX-dev that can generate in $4$ steps instead of 28 steps. However, the model is also more difficult to fine-tune due to distillation (miao2024tuning; jia2024reward), posing challenges for pruning methods that rely on fine-tuning to recover the model performance. Nevertheless, EcoDiff requires less fine-tuning due to its ability to accurately localize redundancy structures. As shown in Table 1, EcoDiff can prune $20\%$ of FLUX-Schnell with only a 0.77 drop in MS COCO FID and a 3.55 drop in Flickr FID, surpassing other baselines. Consequently, EcoDiff can further prune a distilled model to achieve $8.75\times$ speed up compared to the original FLUX-Dev and reduce $20\%$ parameters compared to FLUX-Dev.

### v3

@@ 1 Introduction
Recently, generative models, such as diffusion models and flow matching models, have made remarkable progress in various vision tasks, including text-to-image generation (Brooks et al., 2023; Rombach et al., 2022; Ramesh et al., 2022; Nichol et al., 2022), image inpainting (Saharia et al., 2022a; Lugmayr et al., 2022; Corneanu et al., 2024), super-resolution (Saharia et al., 2022b; Li et al., 2022), and video generation (Ho et al., 2022; Luo et al., 2023; Singer et al., 2022). This progress has been marked by the evolution of generative models toward increasingly larger sizes, transitioning from the U-Net-based Stable Diffusion 1 (SD1) (Rombach et al., 2022) to the larger SDXL (Podell et al., 2024), the transformer-based Stable Diffusion 3 (SD3), and more recently, the FLUX model (Esser et al., 2024; Black Forest Labs, 2024). The most recent FLUX model, with 12 billion parameters, is about 13 times larger than the SD2, which was developed just two years ago (Black Forest Labs, 2024). The rapid growth in size has introduced challenges in application: larger models demand larger GPUs and more computation during inference, limit deployment on smaller compute platforms, and substantially increase the carbon footprint.
@@
Due to the potential challenges induced by larger models, prior methods have explored ways to reduce model size and computation, including model distillation (Meng et al., 2023; Huang et al., 2024) and model pruning (Fang et al., 2023). Compared to standard knowledge distillation that usually trains a student model from scratch, pruning is computationally more efficient. However, pruned models with fewer weights often suffer from degraded generative performance, making retraining necessary to restore their capabilities. Fang et al. (2023) estimate that diffusion model compression can demand 10% to 20% of the original training cost. Therefore, compressing an SD2 can consume up to $40,000$ GPU hours (Stability AI, 2022). Recently, Kim et al. (2024) showed that the cost of pruning SD2 can be reduced to around 300 A100 GPU hours with 0.22M calibration data. For pruning larger models, such as SDXL and FLUX, the retraining burden is more substantial.
@@
One primary reason for the extensive retraining required in prior approaches is due to their coarse pruning criteria. Prior pruning approaches for diffusion models use simple heuristics or one-shot pruning strategies (Kim et al., 2024; Fang et al., 2023), which often fail to balance sparsity and performance. To improve the pruning selection criteria, differentiable masking has been explored in classical pruning research (Louizos et al., 2018; Wen et al., 2016; Liu et al., 2015; Han et al., 2016; Yuan and Lin, 2006) and has recently been adapted to large language models (LLMs) (Fang et al., 2024). However, several unique characteristics of vision generative models make the application of differentiable masking difficult. For instance, vision generative models are Markovian (Song et al., 2021a; b; Lipman et al., 2021), where each generation step only relies on the previous state. Consequently, a minor change at an intermediate generation step can cause a ripple effect that significantly distorts the final result.
@@
While prior works often require extensive retraining to recover model quality after pruning, we show that this step can be made far more lightweight. By combining pruning with targeted post-pruning adaptation, such as low-rank (LoRA) fine-tuning (Hu et al., 2022) or full-model retraining, our approach enables substantial quality recovery without prohibitive compute. This makes pruning practical even for large-scale diffusion models like SDXL, where retraining can otherwise be resource-intensive. However, end-to-end mask learning results in an extremely long gradient chain across all generation steps, leading to significant memory demands. For instance, performing end-to-end pruning on SDXL requires approximately 1400 GB of VRAM–equivalent to the capacity of $15$ NVIDIA H100 GPUs. To address this, we adopt gradient checkpointing (Chen et al., 2016) and introduce time step gradient checkpointing. This approach reduces the VRAM usage for SDXL from 1400 GB to under 30 GB, with a minor overhead of an additional forward pass. An overview of end-to-end pruning and gradient checkpointing is shown in Figure 2.
We show that our method is general and scalable by pruning the most recent diffusion model (SDXL) and flow matching model (FLUX) in Section 6.2. Recent studies have shown that step distillation can greatly speed up inference by reducing generation steps (Salimans and Ho, 2022; Meng et al., 2023). However, step-distilled models are difficult to retrain due to their non-smoothness (Miao et al., 2024; Jia et al., 2024). We show that we can prune step-distilled models in Section 6.2 by pruning FLUX-schnell.
@@ 2 Related Work
Efficient diffusion models: Several works focus on enhancing the efficiency of diffusion models at inference time. DDIM formulates diffusion processes with non-Markovian transformations, reducing the required diffusion steps for high-quality generation (Song et al., 2021a). The Latent Diffusion Model adopts a two-stage generation process, performing the diffusion process in latent space and decoding it back to the image domain using a Variational Autoencoder (VAE) (Rombach et al., 2022; Kingma and Welling, 2014). In addition, some prior works accelerate diffusion inference by sharing intermediate variables to reduce redundant computations (Ma et al., 2024b; a; So et al., 2023). However, these methods are often limited to specific architectures and do not generalize well to newer diffusion models incorporating multi-modal attention. Early stopping methods attempt to terminate the diffusion process once satisfactory generation is reached, but they can sacrifice fine-grained details (Lyu et al., 2022). Distillation-based approaches train a student model that is either smaller or can generate with fewer diffusion steps (Salimans and Ho, 2022; Meng et al., 2023; Gu et al., 2023; Hsiao et al., 2024). Nevertheless, model distillation requires extensive retraining.
@@
Diffusion training efficiency: Beyond inference speedup, research also targets the efficiency of diffusion model training. These efforts range from data-centric acceleration, such as lossless speedup through unbiased dynamic data pruning (Qin et al., 2024) or approaches for data-efficient training via sample reweighting (Li et al., 2025), to architectural and objective acceleration that utilizes masked transformers for faster training (Zheng et al., 2024), patched denoising diffusion models for high-resolution synthesis (Ding et al., 2023), general fast diffusion model designs (Wu et al., 2023), and min-SNR weighting strategies (Hang et al., 2023). These approaches primarily target training-time efficiency, whereas our work focuses on inference-time structural compression and is complementary to the above techniques.
@@
Model pruning: Pruning reduces model size, thereby lowering both memory requirements for loading the model and computation demands during inference. Many classical pruning approaches learn a mask to prune CNN models (Louizos et al., 2018; Wen et al., 2016; Liu et al., 2015; Han et al., 2016; Yuan and Lin, 2006; Feng and Darrell, 2015). Recent advances in model pruning have predominantly focused on large language models (LLMs), where techniques such as unstructured and semi-structured pruning eliminate connections between neurons, and structured pruning targets neurons, attention heads, and layers (Kwon et al., 2022b; Fang et al., 2024; Zhang et al., 2024b; a; Yu et al., 2018; Ma et al., 2023; Kwon et al., 2022a; Sun et al., 2023). In contrast, pruning methods for diffusion models remain relatively underexplored. These approaches vary in their granularity: some involve structural pruning, while others focus on unstructured pruning. Some structural approaches (Castells et al., 2024; Li et al., 2023; Zhao et al., 2024) perform occlusion-based pruning that calculates an importance score for occluding a part through exhaustive search, but its high complexity limits scalability to complex pruning scenarios with many candidates. Unstructured pruning has also been explored, for example, Efficient Pruning of Text-to-Image Models (EPTI) (Ramesh and Zhao, 2024) uses criteria such as magnitude and WANDA, reporting strong results without demanding the extensive retraining common in many other pruning approaches. Fang et al. (2023) uses gradient information as a proxy for neuron importance. Compared to the occlusion method, it scales to more complex pruning cases. Other approaches (Kim et al., 2024) prune certain redundant blocks and recover the model with feature distillation to further reduce the pruning cost. Still, these methods require excessive retraining to retain the model performance.
@@ 3 Preliminaries
Modules in transformer blocks. Transformer blocks are one of the major building blocks in U-Net diffusion models (Rombach et al., 2022; Podell et al., 2024). In addition, recent diffusion transformers employ fully transformer-based architectures. This makes transformer blocks a primary target for pruning to improve efficiency. A transformer block consists of multi-head attention (MHA) and feed-forward network (FFN) (Vaswani et al., 2017). The MHA with $h$ attention heads is defined as,
@@
where $W_{1}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{ff}}}$ and $W_{2}\in\mathbb{R}^{d_{\text{ff}}\times d_{\text{model}}}$. $\sigma(\cdot)$ represents an activation function like GELU or GeGLU (Hendrycks and Gimpel, 2016; Shazeer, 2020).
@@ 5.3 Continuous Relaxation of Discrete Masking
Our pruning objective in Equation 6 involves the calculation of $L_{0}$ norm, which is not differentiable. Hence, we adopt the continuous relaxation of the sparse optimization by applying hard-concrete sampling originally proposed by Louizos et al. (2018). For a scalar continuous masking variable $\hat{\mathcal{M}}\in[0,1]$, the hard-concrete sampling is as follows:
@@ 5.4 Time Step Gradient Checkpointing
To perform the end-to-end pruning shown in Algorithm 1, the backpropagation needs to traverse all generation steps. This necessitates storing all intermediate variables across all steps, leading to a substantial increase in memory usage during mask optimization. We implement time-step gradient checkpointing to address the memory challenge. Traditional gradient checkpointing reduces memory usage by storing selected intermediate outputs and recomputing forward passes between checkpoints (Chen et al., 2016; Gruslys et al., 2016). However, traditional checkpointing only stores intermediate values within a single forward pass, thereby limiting its utility for diffusion models, which require multiple model forward passes across time steps. To address this, we design time step gradient checkpointing, which stores intermediate denoised latent $\hat{z}$ after each denoising step. For a diffusion process with $T$ steps, end-to-end backpropagation has a memory complexity of $O(T)$. By implementing time-step gradient checkpointing, we reduce memory complexity to $O(1)$, making it independent of $T$. The runtime complexity remains $O(T)$, with the only added cost being an additional forward pass to recompute activations as needed. We present Algorithm 2 to show how to apply time step gradient checkpointing to calculate the loss gradient w.r.t. the mask control variable $\bm{\lambda}$.
@@ 6.1 Setup
Model: We prune state-of-the-art (SOTA) models: SDXL, FLUX-dev, and FLUX-schnell (Rombach et al., 2022; Podell et al., 2024; Black Forest Labs, 2024). These models cover diffusion and flow-matching, as well as UNet and DiT architecture. Pruning details: We only use $100$ text prompts from GCC3M (Sharma et al., 2018) for pruning, and a computation budget of 10 A100 GPU hours. Baselines: We compare our method with the SOTA diffusion pruning methods, DiffPruning (Fang et al., 2023), BK-SDM (Kim et al., 2024), and FLUX-Lite (Daniel Verdú, 2024). For BK-SDM, we only run on SD-XL, since the method is designed for U-Net models. We also include the per-step loss variant for mask learning as a baseline. Per-step loss uses identical settings as EcoDiff, except the masking loss is changed to a per-step reconstruction loss. Evaluation metrics: We select Fréchet Inception Distance (FID) (Heusel et al., 2017), CLIP score (Radford et al., 2021), and Structural Similarity Index Measure (SSIM) (Wang et al., 2004). We evaluate under $5,000$ MS COCO and Flickr 30k data samples respectively (Lin et al., 2014; Young et al., 2014). All experiments use a single NVIDIA A100 GPU with 80GB VRAM. More details in Appendix B.
@@ 6.2 Main Results
SDXL is one of the SOTA diffusion models that employs a U-Net architecture with 2.62.6B parameters (Podell et al., 2024). We compare the pruning results with BK-SDM, DiffPruning, and our per-step loss variant. From Table 1 and Figure 3, EcoDiff outperforms baseline methods under the same computation budget. We also observe low SSIM scores for all our pruned models compared to conventional pruning methods (Fang et al., 2023), with all values below 0.65. This suggests that EcoDiff prioritizes semantic integrity (high FID/CLIP) over pixel-level mimicry (low SSIM). The low SSIM is attributed to subtle, perceptually acceptable shifts in generated textures or fine-detail placement, which are heavily penalized by pixel-level metrics but do not detract from the overall image quality or semantic adherence to the prompt. As shown in Figure 3, the semantic fidelity and fine-grained quality of our generated images are noticeably superior to those of the baseline methods. We explain more in Section 6.5 and provide results on SD2 in Appendix D.
@@
Recently, flow-matching has gained attention due to its superior performance. Currently, the best-performing vision generative model is FLUX-Dev, a flow DiT model with 12B parameters (Black Forest Labs, 2024). Due to the DiT architecture, BK-SDM is not applicable to FLUX. Hence, we compare with DiffPruning and FLUX-Lite, a community-pruned model leveraging redundancy of middle layers. According to Table 1, both EcoDiff and FLUX-Lite preserve the model performance. However, the FLUX-Lite 8B model was derived from a large-scale, post-hoc retraining effort, requiring 1120 hours on H200 GPUs to achieve 33% sparsity (Hugging Face Community, 2024), we highlight this comparison to demonstrate that EcoDiff achieves comparable quality and structural sparsity (20%) using only 10 A100 GPU hours for the mask learning stage, positioning our method as a vastly more efficient and accessible alternative to resource-intensive retraining campaigns.
@@
FLUX-Schnell is the step-distilled model from FLUX-dev that can generate in $4$ steps instead of 28 steps. However, the model is also more difficult to fine-tune due to distillation (Miao et al., 2024; Jia et al., 2024), posing challenges for pruning methods that rely on fine-tuning to recover the model performance. Nevertheless, EcoDiff requires less fine-tuning due to its ability to accurately localize redundancy structures. As shown in Table 1, EcoDiff can prune $20\%$ of FLUX-Schnell with only a 0.77 drop in MS COCO FID and a 3.55 drop in Flickr FID, surpassing other baselines. Consequently, EcoDiff can further prune a distilled model to achieve $8.75\times$ speed up compared to the original FLUX-Dev and reduce $20\%$ parameters compared to FLUX-Dev.